OpenStudy (anonymous):

Describe the distributive property? How is it different from other properties? Can you give an example?

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

What is the distributive property? List other types of properties

OpenStudy (anonymous):

@texaschic101

OpenStudy (youngstudier):

Let's see, Symmetric property: If 0.5 = 1/2, then 1/2 = 0.5 Transitive Property: If x = 1/4, and 1/4 = z, then x = z. Commutative Property: of Addition: a + b = b + a, of Multiplication: ab = ba Associative Property: of Addition: (A + b) + c = a + (b + c), of Multiplication: (ab)c = a(bc)

OpenStudy (youngstudier):

Distributive Property: of Addition: a (b+c) = ab + ac Of Subtraction: a (b-c) = ab - ac

OpenStudy (youngstudier):

I'm not sure exactly how to explain how they're different.

OpenStudy (texaschic101):

All the distributive property does is multiply what is outside the parenthesis, with everything inside the parenthesis. Ex. 2(3x + 5) = 2 * 3x + 2 * 5 = 6x + 10

OpenStudy (anonymous):

can you describe what the distributive property is @texaschic101

OpenStudy (texaschic101):

You are basically multiplying the number outside the parenthesis, with everything inside the parenthesis. This gets rid of the parenthesis. a(b + c) = a*b + a*c a(b - c) = a*b + a*-c

OpenStudy (anonymous):

ok and the example?

OpenStudy (texaschic101):

3(x + 4) = 3 * x + 3 * 4 = 3x + 12

OpenStudy (anonymous):

when i tell my teacher the example .. do i just tell him that?

OpenStudy (texaschic101):

if all you have to do is give an example...then yes. But if you have to explain....then just say : Take the number outside of the parenthesis and multiply it by everything inside the parenthesis...and this will get rid of the parenthesis.

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

how is it different from other properties? @texaschic101

OpenStudy (anonymous):

Genny, first tell me in your own words what the distributive property is

OpenStudy (anonymous):

its when you take the numbers outside and multiply them with everything inside:)

OpenStudy (anonymous):

ok, now pick a second property and tell me in your own words what it is and what it does

OpenStudy (anonymous):

the commutative property. . For addition, the rule is "a + b = b + a"; in numbers, this means 2 + 3 = 3 + 2. For multiplication, the rule is "ab = ba"; in numbers, this means 2×3 = 3×2.

OpenStudy (anonymous):

Hum . . . I'm not sure that sound like it is in your own words. What is happening with the commutative property? Look at the first piece a+b and b+a. What happened why are they different but the same?

OpenStudy (anonymous):

its just basically the opposite of each other

OpenStudy (anonymous):

yes, the commutative property means they can be in any order and still result in the same conclusion. How is this different from the distributive one?

OpenStudy (anonymous):

idk :(

OpenStudy (anonymous):

The answer is kinda tricky, but simple. The answer lies in their description. You first describe one and then say, where as the (next one) does (such and such . . .). Let's look at a better comparative example. Tell me if we can use the commutative property with subtraction.

OpenStudy (anonymous):

is 7-2 the same as 2-7?